Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
gherardini - 22:32 Tuesday 25 November 2025 (68252) Print this report
Operator Report - Afternoon shift
The work on HWS-DET vs NE misalignment went on in the afternoon with the ITF locked up to CARM_NULL_1F (#68242), work completed at 19:00UTC; then I relock the cavities and switched off the PR CHRoCC; I left the cavities locked on the infrared. Sub-system reports

TCS

- 19:10UTC: PR CHRoCC VDAC 4.19 --> 0.0 (3600 sec ramp time).


Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×